OpenStudy (anonymous):

Simplify the expressions. Show work. (5x^4 - 3x^3 + 6x)-(3x^3 + 11x^2 -8x) I know there are two ways to do this but I am trying one way and having a problem with one part. So far I have 5x^4 - 3x^3 + 6x - 3x^3 + 11x^2 - 8x (5x^4 - 3x^3) + (3x^3 - 11x^2) + 6x + 8x From here not so sure assuming I have it right so far that is. Does the (5x^4 - 3x^3) become 2x with no ^# or do I subtract them as well that makes 2x^1 which would still be 2x.

Miracrown (miracrown):

You can only subtract like powers. So, 5x^4 - 3x^4 = (5-3)x^4 = 2x^4

Miracrown (miracrown):

But you can't subtract 5x^4-3x^3.

Miracrown (miracrown):

i.e. 5x^4 - 3x^3 is as simple as it could get.

OpenStudy (anonymous):

5x^4 - 3x^3 + 6x - 3x^3 + 11x^2 - 8x there are sign problem here

OpenStudy (anonymous):

So it doesn't change then? It stays 5x^4 - 3x^3? So I have it completely wrong?

OpenStudy (anonymous):

5x^4 - 3x^3 + 6x - 3x^3 -* 11x^2+ 8x

OpenStudy (anonymous):

5x^4 - 3x^3 + 6x - 3x^3 - 11x^2 + 8x

Miracrown (miracrown):

You want to reorder the terms. So that all the x^4's are grouped, x^3's are grouped, etc. Place parentheses around like terms

Miracrown (miracrown):

(5x^4 - 3x^3 + 6x)-(3x^3 + 11x^2 -8x) --> 5x^4 + ( - 3x^3 - 3x^3 ) -11x^2+ (6x + 8x)

Miracrown (miracrown):

Then, factor out the x's from each set of parentheses -- like this: 5x^4 + ( - 3 - 3 )x^3 -11x^2+ (6 + 8)x

Miracrown (miracrown):

Could you do it on a test? Knowing that?

OpenStudy (anonymous):

So, 5x^4 + ( - 3x^3 - 3x^3 ) -11x^2+ (6x + 8x) 5x^4 + 11x^2+ 14x ?

Miracrown (miracrown):

almost, but the x^3 terms don't cancel because they're both negative.

Miracrown (miracrown):

Instead, they combine to make -6x^3.

OpenStudy (anonymous):

Does the + become - then like this? 5x^4 - 6x^3 + 11x^2+ 14x

Miracrown (miracrown):

Yep.

OpenStudy (anonymous):

Great. Thanks so much for helping. Helped a lot.

Miracrown (miracrown):

Yw (: